# Vision-Only Gaussian Splatting for Collaborative Semantic Occupancy Prediction

### v1

@@ Introduction
Datasets. OPV2V is a large-scale collaborative perception dataset collected in the CARLA simulator (Dosovitskiy et al. 2017) using the OpenCDA autonomous driving simulation framework with Vehicle-to-Vehicle (V2V) communication, with each sample containing multi-sensor data from 2-7 vehicles. The original OPV2V does not provide semantic occupancy labels, so we use the enhanced Semantic-OPV2V released by CoHFF (Song et al. 2024), which replays the simulations with additional semantic LiDARs. Following the procedure in CoHFF, we aggregate the multi‑agent ground truth to obtain the collaborative semantic occupancy supervision.
@@
Implementation details. Following CoHFF (Song et al. 2024), we utilize a $40\times 40\times 3.2$ meter detection area with a grid size of $100\times 100\times 8$, resulting in a voxel size of $0.4m^{3}$. Unless otherwise noted, each agent default utilizes $|\mathcal{G}|=P=25600$ Gaussians as scene representation, and neighborhood $\rho$ is set to $0.4m$ with attention pooling. We allow CAVs to transmit and share Gaussian primitives for cross-agent fusion. Our experiment incorporates the analysis of 12 semantic labels plus an additional empty label. For optimization, we use AdamW (Loshchilov and Hutter 2017) with weight decay $0.01$. The learning rate warms up for the first 500 iterations to $2\times 10^{-4}$ and then follows a cosine decay. We train for 30 epochs with batch size 8 on a single NVIDIA A100 (40GB). Unless noted otherwise, all components are trained end to end. Following GaussianFormer (Huang et al. 2024b), we use voxel‑wise cross‑entropy loss together with Lovász‑Softmax (Berman, Triki, and Blaschko 2018) loss, which directly targets IoU improvement. The total loss is:
$L\;=\;L_{\mathrm{CE}}\;+\;L_{\mathrm{Lovasz}}\enspace.$ (21)
@@
Following (Cao and De Charette 2022; Huang et al. 2023b; Song et al. 2024), we adopt mean Intersection-over-Union (mIoU) and Intersection-over-Union (IoU) to evaluate the performance of semantic occupancy prediction. For evaluations in subsequent applications, following prior work (Song et al. 2024), we report BEV segmentation using 2D IoU by projecting predicted and ground-truth semantic voxels onto the BEV plane and compute IoU for vehicles, roads and other general objects.

### v2

@@ Introduction
Datasets.
OPV2V is a large-scale collaborative perception dataset collected in the CARLA simulator (dosovitskiy2017carla) using the OpenCDA autonomous driving simulation framework with Vehicle-to-Vehicle (V2V) communication, with each sample containing multi-sensor data from 2-7 vehicles. The original OPV2V does not provide semantic occupancy labels, so we use the enhanced Semantic-OPV2V released by CoHFF (cohff), which replays the simulations with additional semantic LiDARs. Following the procedure in CoHFF, we aggregate the multi‑agent ground truth to obtain the collaborative semantic occupancy supervision.
@@
Implementation Details.
Following CoHFF (cohff), we utilize a $40\times 40\times 3.2$ meter detection area with a grid size of $100\times 100\times 8$, resulting in a voxel size of $0.4m^{3}$. Unless otherwise noted, each agent default utilizes $|\mathcal{G}|=P=25600$ Gaussians as scene representation, and neighborhood $\rho$ is set to $0.4m$ with attention pooling. We allow CAVs to transmit and share Gaussian primitives for cross-agent fusion. Our experiment incorporates the analysis of 12 semantic labels plus an additional empty label. For optimization, we use AdamW (adamw) with weight decay $0.01$. The learning rate warms up for the first 500 iterations to $2\times 10^{-4}$ and then follows a cosine decay. We train for 30 epochs with batch size 8 on a single NVIDIA A100 (40GB). Unless noted otherwise, all components are trained end to end. Following GaussianFormer (huang2024gaussianformer), we use voxel‑wise cross‑entropy loss together with Lovász‑Softmax (berman2018lovasz) loss, which directly targets IoU improvement. The total loss is:
$L\;=\;L_{\mathrm{CE}}\;+\;L_{\mathrm{Lovasz}}\kern 5.0pt.$ (21)
@@
Following (monoscene; tpvformer; cohff), we adopt mean Intersection-over-Union (mIoU) and Intersection-over-Union (IoU) to evaluate the performance of semantic occupancy prediction. For evaluations in subsequent applications, following prior work (cohff), we report BEV segmentation using 2D IoU by projecting predicted and ground-truth semantic voxels onto the BEV plane and compute IoU for vehicles, roads and other general objects.
@@ Conclusion
Acknowledgments
This material is based in part upon work supported by the National Science Foundation under the Center CHORUS with Grant Number CNS-2333487, Army Research Lab under Contract number W911NF-20-2-0026, and gift funding from Amazon AWS. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.